OpenStudy (unimatix):

Prove that both trapezoid rule and Simpson's rule give an estimate within 0.005 of true value of integral?

OpenStudy (unimatix):

Using Maple I've found: trapezoid rule gives 2.432066146 Simpson's rule gives 2.430797145

OpenStudy (unimatix):

\[\int\limits_{a}^{b}\frac{ x^2 }{ \sqrt{(1-x^2)(k^2-x^2)} }\] a = 3 b = 5 k = 2

OpenStudy (unimatix):

@SolomonZelman @pooja195

OpenStudy (unimatix):

Even if someone just has an idea that might be helpful.

OpenStudy (dan815):

error bound on simpsons rule

OpenStudy (anonymous):

you want to confirm this numerically or you want to prove it using the error bound theorem

OpenStudy (dan815):

i think to find error bounds for trapezoidal one

OpenStudy (dan815):

take all right end points and all left end points

OpenStudy (dan815):

the true value is between those 2 answers

OpenStudy (dan815):

if you show that the difference between trap right end points trap left end points method is less than 0.005 then your answer is within 0.005 of the true value

OpenStudy (unimatix):

Okay thanks I'll try and see if I can figure out how to do that.

OpenStudy (dan815):

I am not completely sure about this argument though

OpenStudy (dan815):

it could be possible that both the right end points and the left end points are both greater than the true value, or both less than the true value

OpenStudy (dan815):

yeah throw that method out of the window, there is a more mathematical way

OpenStudy (dan815):

https://www.youtube.com/watch?v=qVXIU6mKank This tells you about simpsons error bound

OpenStudy (unimatix):

Hah that's funny I just found that. I don't see one for the Trapezoid rule though :-/

OpenStudy (dan815):

isnt simpsons just a higher order trap

OpenStudy (unimatix):

I don't really know what that means.

OpenStudy (dan815):

simpsons rule is just like beight a bit smarter about the trapezoidal rule

OpenStudy (dan815):

it will take the average of the areas of the trapezoids in a smarter way

OpenStudy (unimatix):

Okay. So it will be found in a similar way?

OpenStudy (dan815):

yeah u should understand the pictures of these operations that will make it a lot clearer

OpenStudy (dan815):

https://www.youtube.com/watch?v=uc4xJsi99bk

OpenStudy (dan815):

ok i see now

OpenStudy (dan815):

basically for trapezoidal

OpenStudy (dan815):